JET SPECTROSCOPY OF THE 2ν_{8} BAND OF HYDROGEN FLUORIDE TRIMER IN THE 750 CM^{−1} REGION

\underline{THOMAS A. BLAKE} and STEVEN W. SHARPE, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, 3020 Q Avenue, Richland, WA 99352.

Theory predicts a planar ring, oblate symmetric top structure (C_{3h}) for hydrogen fluoride trimer. This ground state structure has been confirmed experimentally for deuterium fluoride trimer by Suhm, et al.^{a} and for hydrogen fluoride trimer by Michael and Lisy.^{b} Both of these experiments involved examining intramolecular stretching modes in the cluster. We have recorded what we believe to be the first rotationally resolved spectra of an intermolecular mode: the first overtone of the ν_{8} (e′′) out-of-plane torsion, which was predicted by Kolenbrander, et al.^{c} This spectrum was produced by expanding a 4% HF-in-helium mixture through a pulsed, slit-jet source and interrogating the jet with the output of a lead salt diode laser spectrometer. A pressure and concentration study showed that most transitions (FWHM 45 MHz) in the 750 cm^{−1} region belong to the same oligimer. Groupings of intense lines are spaced by about 0.284 cm^{−1} suggesting a B value of 0.142 cm^{−1} which is close to the predicted value of 0.13 cm^{−1}. The appearance of the spectrum is, however, far more complicated than that predicted for a perpendicular band of a semi-rigid oblate symmetric top.

aMartin A. Suhm, John T. Farell, Jr., Stephen H. Ashworth, David J. Nesbitt, J. Chem. Phys. 98, 5985 (1993).

bDaniel W. Michael, James M. Lisy, J. Chem. Phys. 85, 2528 (1986).

cKirk D. Kolenbrander, Clifford E. Dykstra, James M. Lisy, J. Chem. Phys. 88, 5995 (1988).